Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4c)
amagazzu - 23:05 Friday 01 August 2025 (67413) Print this report
Operator Report - Afternoon shift

ITF found in DOWN state and in EARTHQUAKE Mode (M 5.9, 11:57 UTC, 164 km SE of Vilyuchinsk, Russia). From 13:24 UTC it was possible to relock, ITF back in Science mode from 14:07 UTC.
At 17:19 UTC the temperature of the Neovan head suddenly increased from 28 to 36 degrees, triggering its safety. To restore the system De Rossi went to the Laser lab and from 18:00 UTC it was possible to relock again.
ITF back in Science Mode from 18:46 UTC, it unlocked again at 19:17 UTC due to the influence of an earthquake (M 6.1, 18:20 UTC, 219 km ESE of Severo-Kuril’sk, Russia). 
Back in Science Mode from 20:07 UTC.
ITF left locked.

Guard Tour (UTC)
18:05 - 18:42
20:57 -

Images attached to this report
Comments to this report:
derossi - 11:10 Saturday 02 August 2025 (67418) Print this report

The chiller was givinig "flow error". Indeed it seems that the flow stopped at first (something blocking it?), and then the temperature started to increase (plot 1). I just stopped and  restarted a few times the chiller, and unplugged both the outlet and inlet pipes, in order to verify that the water was flowing. I did not try to reverse the flow because the chiller restarted before doing this last test. 

The system seems to be returned at its nominal value (plot 2).

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×